OpenStudy (spicytaco1):

will fan A regular pentagon has sides measuring 8 cm. The pentagon is enlarged using a scale factor of . Part a) What are the lengths of each side of the enlarged pentagon? Part b) What is the perimeter of the enlarged pentagon?

OpenStudy (anonymous):

you have to know the scale factor?

OpenStudy (anonymous):

The pentagon is enlarged using a scale factor of 5/2 " translation: each side or dimension is multiplied by 5/2 5/2 = 2.5

OpenStudy (anonymous):

so let's say each side is 6 cm long you multiply each side by 5/2 or 2.5 to get 6*(5/2) = 6*(2.5) = 15 each side is now 15 cm long. This is not the answer as it's just an example

OpenStudy (anonymous):

each side is 20 cm there are 5 sides (this is a pentagon) each side is congruent or equal to one another (it's a regular polygon)